OpenStudy (anonymous):

[02.01]What did Schrodinger's quantum mechanical model eventually show?

OpenStudy (anonymous):

atomic mass Avogadro's constant modern atomic theory Bohr's Theory of the Atom

OpenStudy (anonymous):

helppppp please

OpenStudy (chmvijay):

Bohr's Theory of the Atom

OpenStudy (anonymous):

According to Schrodinger's work with electrons and their orbital paths, what can be determined about the electrons of an atom?

OpenStudy (anonymous):

Most atoms are smaller than electrons. Most atoms contain electrons in the nucleus. Most of the mass of the atom is in electrons. Most of the volume of the atom is occupied by electrons.

OpenStudy (anonymous):

please help

OpenStudy (chmvijay):

Most of the volume of the atom is occupied by electrons.

OpenStudy (anonymous):

]Magnesium has three naturally occurring isotopes. The most common, magnesium-24, accounts for nearly 80% of the magnesium found in nature. The atomic numbers and mass numbers for four different atoms are shown in the table below. Which of these atoms could be an isotope of magnesium?

OpenStudy (anonymous):

If a chemist has 2.3 moles of arsenic that has a molar mass of 74.92 g/mol, which would be the best way to find out how many atoms of arsenic are in the sample?

OpenStudy (anonymous):

Multiply 2.3 by the molar mass of arsenic. Divide 2.3 by the molar mass of arsenic. Multiply 2.3 by Avogadro's number. Divide 2.3 by Avogadro's number.

OpenStudy (chmvijay):

Multiply 2.3 by Avogadro's number

OpenStudy (anonymous):

Silicon has three naturally occurring isotopes. Silicon-28 has a mass of 27.98 and a relative abundance of 92.23%. Silicon-29 has a mass of 28.98 and a relative abundance of 4.68%. Silicon-30 has a mass of 29.97 and a relative abundance of 3.09%. What is the weighted average atomic mass of silicon?

OpenStudy (anonymous):

27.99 g/mol 28.09 g/mol 28.98 g/mol 29.00 g/mol

OpenStudy (anonymous):

hhhheeeeeeeeeeeeeeellllllllpppppppppppppppp

OpenStudy (chmvijay):

IDK!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!:(

OpenStudy (anonymous):

If a sample of pure iodine contains 4.69 x 1022 atoms of iodine, how many moles of iodine are in the sample?

OpenStudy (anonymous):

0.08 mol 0.13 mol 7.78 mol 12.84 mol

OpenStudy (anonymous):

helppppppp

OpenStudy (chmvijay):

6.023*10^23 iodine atoms = 1mole 4.69*10 ^22iodine atoms =1*4.69*10 ^22 / 6.023*10^23 u calculate LOL:)